The following returns have been estimated for Security T and Security S:

Scenario

Security T

Security S

1

20%

10%

2

13%

-6%

3

15%

20%

Each scenario is equally likely to occur, and you plan to invest 70% in Security T and 30% in Security S. What is the expected return of the portfolio? Round your answer to the nearest tenth of a percent.
Answer #1
Scenario Security T Security S
1 20% 10%
2 13% -6%
3 15% 20%

Each scenario is equally likely to occur means that the probability of occurrence of each scenario is 1/3

Probability of scenario 1 = p1 = 1/3

Probability of scenario 2 = p2 = 1/3

Probability of scenario 3 = p3 = 1/3

Expected return when the probability of occurrence of different scenarios is given is calculated using the formula:

Expected return = E[R] = p1*R1 + p2*R2 + p3*R3

Expected return on Security T = E[RT] = (1/3)*20% + (1/3)*13% + (1/3)*15% = 16%

Expected return on Security S = E[RS] = (1/3)*10% + (1/3)*(-6%) + (1/3)*20% = 8%

Portforlio

Weight of Security T in portfolio = WT = 70%, Expected return on Security T = E[RT] = 16%

Weight of securty S in portfolio = WS = 30%, Expected return on Security S = E[RS] = 8%

Expected return of the portfolio is calculated using the formula:

Expected return of portfolio = E[RP] = WT*E[RT] + WS*E[RS] = 70%*16% + 30%*8% = 13.6%

Answer -> Expected return of portfolio = 13.6%